OpenStudy (anonymous):

Use quadratic formula to solve the equation -x^2= 16x+46

OpenStudy (jdoe0001):

\(\bf -x^2= 16x+46\implies -1x^2-16x-46=0\implies 1x^2+16x+46=0 \\ \quad \\ \textit{quadratic formula}\\ \begin{array}{llll} y=&ax^2+&bx+&c\\ &\uparrow&\ \uparrow&\uparrow\\ &a&\ b&c\\ &1x^2&+16x&+46 \end{array} x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

How did you get positive 16 and 46

OpenStudy (jdoe0001):

well, you can use the negative version, just the same :), it''d be just a bit cleaner using the positive one, but \(\bf -x^2= 16x+46\implies -1x^2-16x-46=0 \quad \times -1\textit{ both sides}\\ \quad \\ \implies 1x^2+16x+46=0\)

OpenStudy (jdoe0001):

recall that -1 * 0 = 0

OpenStudy (whpalmer4):

\[-x^2=16x+46\]Rearrange to the standard \[ax^2+bx+c=0\]by adding \(x^2\) to each side:\[-x^2+x^2=x^2+16x+46\]\[0=x^2+16x+46\]\[a=1,\,b=16,\,c=46\]and you're off to the races